# EDEN: Efficient Dual-Layer Exploration Planning for Fast UAV Autonomous Exploration in Large 3-D Environments

### v1

@@ V-A Implementation Details
$\ddot{\psi}_{m}=1.57\,rad/s^{2}$ $\lambda_{a}=0.6$ $\lambda_{arc}=1.0$ $\lambda_{e}=0.3$
$\lambda_{p}=4000$ $\lambda_{d}=350$ $\lambda_{t}=700$
V-B Simulation Results
In this section, we compare the proposed method against [20, 16] in four simulation environments. Each method runs 10 times in a simulation environment. We record the exploration time, UAV speed, traveled distance, and computational cost for evaluation (shown in Table. II).
  • FALCON [20]: This method plans the exploration sequence of unexplored regions by solving ATSP. After planning the exploration sequence, a time-optimal B-spline trajectory is generated utilizing [38].
  • FUEL [16]: This method plans the exploration sequence of frontiers through solving ATSP. It also uses time-optimal B-spline trajectories for exploration.
@@ V-B Simulation Results
(a) Proposed.
(b) FALCON.
(c) FUEL.
(d) Explored Volume.
(e) Proposed.
(f) FALCON.
(g) FUEL.
(h) Explored Volume.
(i) Velocity Color.
(j) Proposed.
(k) FALCON.
(l) FUEL.
(m) Explored Volume.
Figure 4: The mapping results and the executed trajectory of all the methods. The color of the trajectory denotes the velocity of the UAV. Each row illustrates the simulations in one environment. The right column illustrates the explored volume vs. time.
@@
The first simulation environment is the DARPA Tunnel with a scale of $42\times 20\times 2\,m^{3}$. Each method runs $75\,s$. This environment is relatively simple, so the main factor that affects the exploration efficiency is the speed of the UAV. As shown in Fig.4 (a-c), the proposed method maintains a higher speed than other methods ($10.4\%$ faster than FALCON and $28.6\%$ faster than FUEL). As a result, the proposed method is $5.2\%$ more efficient than FALCON and $16.1\%$ more efficient than FUEL.
@@
The second simulation environment is the Classical Office with a scale of $15\times 30\times 2\,m^{3}$. Each method runs $130\,s$. Compared with DARPA Tunnel, Classical Office is more complex. The walls force the UAV to keep turning which slows down the speed of the UAV. Thanks to the curvature-penalized viewpoints planning and ASEO trajectory, the proposed method is able to avoid sharp turns and generate high-speed trajectories as shown in Fig. 4 (e-g). The UAV speed of the proposed method is $12.6\%$ faster than FALCON’s and $36.6\%$ faster than FUEL’s. Since this environment is more open than DARPA Tunnel, the computational costs grow. The computational cost proposed method is $28.9\%$ lower than FALCON’s and $36.6\%$ lower than FUEL’s. As a result, the proposed method is $12.1\%$ more efficient than FALCON and $40.4\%$ more efficient than FUEL.
The third simulation environment is the Maze with a scale of $40\times 40\times 2\,m^{3}$. Each method runs $500\,s$. As shown by Fig. 4 (j-l), the proposed method keeps high speed in this environment. Because this environment is larger than previous ones, the computational costs of FALCON and FUEL are much higher. The computational cost of the proposed method also grows, but only $12.1\%$ compared to simulation in Classical Office. This is because EOCH can solve ATSP efficiently. The path lengths of the proposed method and FALCON remain similar, which are shorter than FUEL’s, though FALCON and FUEL have better solutions for ATSP. This is because their implicit optimal assumptions fail most of the time. Thus, finding a perfect ATSP solution in exploration is usually not cost-effective. The results show the proposed method’s exploration efficiency is $14.0\%$ more efficient than FALCON and $42.6\%$ more efficient than FUEL.
@@
The last simulation environment is the Large Tunnel with a scale of $251.5\times 327.5\times 2\,m^{3}$. Since this environment is very large, we adopt $v_{m}=4.0m/s$ and $a_{m}=6.0m/s^{2}$ to save the simulation time. The purpose of this trial is to test the computational cost of the proposed method in an extremely large environment. The comparative method FALCON and FUEL fail to run in this environment due to their memory usage exceeding the maximum 16G memory of the computer. On the contrary, the proposed method can accomplish the exploration as shown by Fig. 5. The average computational cost of planning is $20.48\,ms$ which is low enough to plan in real time. The average speed is $3.32\,m/s$. The average exploration time is $2441.32\,s$. The result shows the proposed method is able to explore large-scale environments.
@@
(b) Executed Trajectory and Mapping Result.
(c) Outdoor Environment.
(d) Exploration Volume.
Figure 6: Experiment results of real-world exploration.
V-C Real-World Experiment
To evaluate the proposed method further, we deploy our method on a self-made UAV (Fig. 6(a)) to explore a $42\times 21\times 3\,m^{3}$ outdoor environment (Fig. 6(c)). The exploration results are shown in Fig. 6 (b) and (d). The exploration cost $145.1\,s$ to travel $260.9\,m$, explore $2520\,m^{3}$ of the environment is explored. During the exploration, the average velocity of the UAV is $1.79m/s$, and the average planning time is $18.5\,ms$.
The simulation results show that the proposed method can plan high-speed exploration trajectories in real-time even in large 3-D environments. The real-world experiments show that the proposed method is also effective.
VI conclusion
In this work, we propose an efficient dual-layer exploration planning method (EDEN). By finding long-term routings through the proposed EOHC algorithm, deciding short-term targets through curvature-penalized evaluation, and generating ASEO trajectories, the proposed method is able to explore large environments efficiently with low computational cost and high speed. The simulation indicates the exploration efficiency of the proposed method is $5.2\%-14.0\%$ higher than state-of-the-art methods. The computational cost of the proposed method is low enough to enable exploration in large-scale environments.

### v2

@@ VI-A Implementation Details
$\ddot{\psi}_{m}=1.57\,rad/s^{2}$ $\lambda_{a}=0.6$ $\lambda_{e}=0.3$ $\lambda_{p}=4000$
$\lambda_{d}=350$ $\lambda_{t}=700$ $r_{u}=0.35\,m$ $d_{th}=0.6\,m$
$\psi_{th}=0.25\,rad$ $t_{r}=2.0\,s$ $t_{o}=0.1\,s$
(a) Proposed.
(b) FALCON.
(c) FUEL.
(d) Proposed.
(e) FALCON.
(f) FUEL.
(g) Proposed.
(h) FALCON.
(i) FUEL.
(j) Velocity.
(k) Proposed.
(l) FALCON.
(m) FUEL.
(n) Explored Volume (DARPA Tunnel).
(o) Explored Volume (Classical Office).
(p) Explored Volume (Large Maze).
(q) Explored Volume (City).
Figure 4: The mapping results and the executed trajectory of all the methods. The color of the trajectory (j) denotes the velocity of the UAV. Each row illustrates the simulations in one environment. The last row illustrates the explored volume vs. time.
@@
VI-B Simulation Results
In this section, we compare the proposed method against [22, 19] in four simulation environments. For fair comparison, we modify the velocity and acceleration optimizations of the comparative methods. The modification constrains their trajectory generation within the maximum linear velocity and acceleration, rather than within the three axes individually. Except for the size of the exploration space, the other parameters of the comparative methods are not modified. Each method runs 10 times in a simulation environment. We record the exploration time, UAV speed, traveled distance, and computational cost for evaluation (shown in Table. II).
  • FALCON [22]: This method plans the exploration sequence of unexplored regions by solving coverage-ATSP. After planning the exploration sequence, a time-optimal B-spline trajectory is generated utilizing [44].
  • FUEL [19]: This method plans the exploration sequence of frontiers through solving frontier-ATSP. It also uses time-optimal B-spline trajectories for exploration.
@@ VI-B Simulation Results
The first simulation environment is the DARPA Tunnel with a scale of $42\times 20\times 2\,m^{3}$. Each method runs 75s75\,s. This environment is relatively simple, so the main factor that affects the exploration efficiency is the speed of the UAV. As shown in Fig.4 (a-c), the proposed method maintains a higher speed than other methods ($10.4\%$ faster than FALCON and $28.6\%$ faster than FUEL). As a result, the proposed method is $5.2\%$ more efficient than FALCON and $16.1\%$ more efficient than FUEL. Fig. 4(n) compares the explored volume of the three methods.
@@
The second simulation environment is the Classical Office with a scale of $15\times 30\times 2\,m^{3}$. Each method runs 130s130\,s. Compared with DARPA Tunnel, Classical Office is more complex. The walls force the UAV to keep turning, which slows down the speed of the UAV. Thanks to the curvature-penalized viewpoints planning and ASEO trajectory, the proposed method is able to reduce sharp turns and generate high-speed trajectories, while FALCON usually decelerates when meeting walls, and FUEL can reach high speed only in explored space, as shown in Fig. 4(d-f). This is because the B-spline trajectories of FALCON and FUEL only try to minimize the current trajectory duration without considering continuous exploration. Moreover, the exploration efficiencies of FALCON and FUEL are limited because their implicit optimal assumptions are usually not fulfilled due to the wall obstacles and open space of this environment. The UAV speed of the proposed method is $12.6\%$ faster than FALCON’s and $36.6\%$ faster than FUEL’s. Since this environment is more open than DARPA Tunnel, the computational costs of the three methods grow. The computational cost of the proposed method is $29.3\%$ lower than FALCON’s and $37.7\%$ lower than FUEL’s. The traveled distances of the three methods are similar. As a result, the proposed method is $12.1\%$ more efficient than FALCON and $40.4\%$ more efficient than FUEL. Fig. 4(o) compares the explored volume of the three methods.
The third simulation environment is the Maze with a scale of $40\times 40\times 2\,m^{3}$. Each method runs 500s500\,s. As shown by Fig. 4(g-i), the proposed method keeps a high speed in this environment. Because this environment is larger than previous ones, FALCON and FUEL have to spend more computing time on solving larger-scale ATSP. The computational cost of the proposed method also grows, but only $12.1\%$ compared to simulation in Classical Office. This is because EOHDT can solve ATSP efficiently. Similar to the result in Classical Office, the traveled distance of the proposed method is slightly longer than FALCON’s but shorter than FUEL’s. This is caused by two factors. First, the curvature-penalized viewpoint planning makes the proposed method tend to choose to explore the target that can keep high UAV speed, which may leave some isolated unexplored corners and lead to the back-and-forth movements. Second, the routing solved by the proposed EOHDT is sub-optimal. Even so, the sacrifice of the traveled distance trade offers a lower computational cost and a faster UAV speed.
The fourth simulation environment is the City with a scale of $40\times 40\times 8.5\,m^{3}$. Each method runs 1200s1200\,s. Compared with previous environments, City has a larger size and contains denser obstacles, which pose significant challenges to the computational efficiency of each method. As shown by Fig. 4(k-m), the proposed method executes a straighter trajectory and keeps a higher speed, while comparative methods’ trajectories are more curved and their speeds are lower. The average computation time for both FALCON and FUEL exceeds 0.5s0.5\,s, which has a significant impact on the exploration of real-time planning. In contrast, thanks to EOHDT, the proposed method is able to plan in real-time with massive boundary regions. Consequently, the proposed method is $89.7\%$ more efficient than FALCON and $93.5\%$ more efficient than FUEL. Moreover, the traveled distance of the proposed method is shorter than that of FALCON. This is because City contains dense building obstacles, which are far from FALCON’s implicit optimal assumption.
@@
The last simulation environment is the Large Tunnel with a scale of $251.5\times 327.5\times 2\,m^{3}$. Each method runs 3000s3000\,s. Since this environment is very large, we adopt $v_{m}=4.0m/s$ and $a_{m}=6.0m/s^{2}$ to save the simulation time. The purpose of this trial is to test the computational cost of the proposed method in an extremely large environment. The comparative methods, FALCON and FUEL, fail to run in this environment due to their high memory usage and large computation. On the contrary, the proposed method can accomplish the exploration as shown by Fig. 5. The average computational cost of planning is 20.48ms20.48\,ms, which is low enough to plan in real time. The average speed of the UAV is 3.32m/s3.32\,m/s. The explored volume of the proposed method is shown in Fig. 5(b). The average exploration time is 2441.32s2441.32\,s. The result shows the proposed method is able to explore large-scale environments.
@@
(b) Courtyard.
(c) Park.
(d) Mapping Result (Courtyard).
(e) Mapping Result (Park).
(f) Explored Volume (Courtyard).
(g) Explored Volume (Park).
Figure 6: Experiment results of real-world explorations.
VI-C Real-World Experiment
To evaluate the proposed method further, we deploy our method on a self-made UAV (Fig. 6(a)), and test it in two outdoor environments. The first environment is the $Courtyard$ with a scale of $42\times 21\times 3\,m^{3}$ as shown by Fig. 6(b). This environment is composed of trees and bushes. The exploration results are shown in Fig. 6(d) and Fig. 6(f). The exploration takes 145.1s145.1\,s to travel 260.9m260.9\,m, explore $2520\,m^{3}$ of the environment. During the exploration, the average velocity of the UAV is 1.79m/s1.79\,m/s, and the average planning time is 18.5ms18.5\,ms. The second environment is the $Park$ with a scale of $30\times 25\times 3\,m^{3}$ as shown by Fig. 6(c). This environment is composed of a pergola, uneven slopes, and cluttered trees, which is more challenging than the first environment. The exploration results are shown in Fig. 6(e) and Fig. 6(g). The exploration takes 145.5s145.5\,s to travel 256.4m256.4\,m, explore $2131\,m^{3}$ of the environment. The average velocity of the UAV is 1.77m/s1.77\,m/s, and the average planning time is 14.1ms14.1\,ms. Since this environment is more cluttered and smaller, the number of viewpoints to be evaluated and the number of nodes in ATSP are fewer, which reduces the UAV speed and the planning time. The results show the proposed method is able to plan a high-speed trajectory with limited computation resources online.
The simulation results show that the proposed method can plan high-speed exploration trajectories in real-time, even in large 3-D environments. The real-world experiments show that the proposed method is also effective.
VII CONCLUSION
In this work, we propose an efficient dual-layer exploration planning method (EDEN). By integrating the EOHDT algorithm for long-term region routing, the curvature-penalized viewpoint selection for short-term target planning, and the ASEO trajectory generation for high-speed continuous flight, EDEN significantly improves exploration performance. Experimental results show that EDEN achieves up to $89.7\%$ higher exploration efficiency, $10.4\%-69.7\%$ faster UAV speed, and over $28.9\%$ reduction in computational cost compared to state-of-the-art methods. These advantages were consistently observed across multiple complex simulation environments and validated in real-world outdoor tests. The method enables real-time planning in large-scale environments while maintaining high-speed flight and safety. One drawback of the proposed method is that continuity-enhanced exploration planning increases the traveled distance. We plan to prioritize the exploration of isolated unknown regions to reduce redundant movements.